# ExaGRyPE: Numerical General Relativity Solvers Based upon the Hyperbolic PDEs Solver Engine ExaHyPE

### v2

@@ 4.2.1 Solver Configuration
Once we have a reasonably fine initial mesh, ExaHyPE starts to query the solver objects if a certain area should be refined. In this stage, the solvers do not yet have a solution at hand. Either we hard-code the initial refinement pattern, or we make the algorithm run through a decision logic which analyses “if we had a solution here, would we refine”. This implements a feature-based initial refinement. The present case studies all use hard-coded patterns, which we either write into the C++ template or inject the implementation of the C++ code directly from Python via my_solver.set_implementation(refinement_criterion
@@ 5.1 Gauge Wave
$\hat{t}$ $=$ $t-\frac{A}{2k\pi}\cos[k\pi(x-t)],$
$\hat{x}$ $=$ $x+\frac{A}{2k\pi}\cos[k\pi(x-t)],\qquad\text{and}$
$\hat{y}$ $=$ $y,\ \hat{z}=z,$
@@
$\alpha=\sqrt{H},~{}\beta^{i}=0,~{}\phi=H^{-1/6},\quad\text{and}$
$K_{xx}$ $=$ $-\frac{\partial_{t}H}{2\alpha}$
@@
$K_{ij,{\rm others}}$ $=$ $0,$
$K=\gamma^{ij}K_{ij}$ $=$ $\frac{K_{xx}}{H}$
@@
The simulation test we report in this subsection utilizes a computational domain $\Omega=[-0.5,0.5]^{3}$ on a regular grid subject to periodic boundary conditions. Periodic boundary conditions are also a newly added feature released with ExaHyPE 2 as opposed to the first-generation code. We divide the domain into 162 volumes per dimension, thus leading to a volume size of $0.006173$. AMR is not enabled for this test.
@@
We illustrate the simulation result at code time $T=0.5$ in Figure 8, where the upper panel gives the profiles of $\tilde{\gamma}_{11}$ along the $x$-axis for different kernels, all employed the first-order formulation: FV, FD4-RK1, FD4-RK2 and FD4-RK4, also with a theoretical solution as a reference; The lower panel gives the $L_{2}$ error along the same line for each solver, calculated as
$L_{2}:=\sqrt{\frac{\sum_{i\in n}(Q^{i}_{\rm simulated}-Q_{\rm Theoretical})^{2}}{n}}.$ (9)
@@
Here $n$ is the total number of sample points. All the solvers yield the correct traveling speed and phase of the wave. However, the Finite Volume solver suffers from severe amplitude damping, and a similar dissipation is also observed in the black hole benchmarks. Consequently, we have chosen to use the FD4 scheme as our default solver.
@@
Figure 8: Upper panel: The profile of $\tilde{\gamma}_{11}$ along $x$-axis at code time $0.5$ for different solvers: FV-RK1 (red solid), FD4-RK1 (blue solid), FD4-RK2 (green dashed) and FD4-RK4 (orange dotted). Lower panel: The corresponding $L_{2}$ error for each solver. The FV solver suffers from a strong dissipation, thus yielding a high error.
On the other hand, different orders of the RK schemes give rather consistent results against the analytical solution. However, their improvement on the long-term evolution of the system can be significant. In our further experiments, a simple Euler integrator showed moderate deviation from the analytical solution at a code time of $T=10$ (not shown). A higher-order temporal scheme could be employed to address this issue. It can be seen more clearly in Figure 9: we show the violations of the Hamiltonian constraints $H$ and the $x$-component of the momentum constraints, $M^{1}$, along the $x$-axis for each integrator in the upper panel. The corresponding $L_{2}$ errors of the violations are presented in the lower panel. Higher-order RK schemes yield significantly lower constraint violations compared to the Euler integrator, indicating superior accuracy.
Figure 9: Upper panel: The profile of violations of $H$ and $M^{1}$ along the $x$-axis at code time $0.5$ for different orders of the RK scheme (1,2 and 4), with an FD4 kernel. Lower panel: The corresponding $L_{2}$ errors for different scheme. A high-order RK time integrator significantly suppresses the constraint violation, indicating better accuracy.
Our tests demonstrate that our code can handle the gauge wave scenario properly using various solvers and timestepping schemes. Due to the time-consuming nature of long-term gauge wave tests, we have not included them in this paper. Similarly, the black hole simulations in the subsequent section do not employ higher-order RK schemes because of their computational intensity. We recognize the importance of these aspects and plan to focus on performance optimization in future work. This will allow us to revisit the long-term behavior of the gauge wave and evaluate the impact of higher-order RK schemes with a more efficient code base of ExaGRyPE.
@@ 5.2 Single Schwarzschild Black Hole
We next examine a single Schwarzschild black hole. This is yet another stationary test, where all numerical ingredients have to be carefully balanced against each other such that the initial setup effectively yields a long-term stable, time-invariant solution. The black hole is placed at the origin of the three-dimensional domain, with an ADM mass $M=1$ and zero spin $S=0$. All initial conditions for this test stem from the TwoPuncture module of the Einsteintoolkit library and use the average initial lapse
$\alpha=\frac{1}{2}\left(\frac{1-\frac{1}{2}\left(M/r\right)}{1+\frac{1}{2}\left(M/r\right)}+1\right).$
Other initial profiles and lapses are used in other numerical codes (cmp. GRChombo). However, a convergence from different setup has been observed. The initial shift vector is set to zero. Also, the extrinsic curvature $K_{ij}$ vanishes in the initial condition: its trace is zero as we assume the maximal slicing $K=0$, and its residual is zero as no linear or angular momentum is presented in this system. Finally, we also assume conformal flatness in the initial condition, i.e. $\tilde{\gamma}_{ij}=1$.
@@
$\phi\equiv\psi^{-2}=\left(1+\frac{M}{2r}\right)^{-2}.$
The domain in this test has a size of $[-9M,9M]^{3}$. We use three levels of AMR based on the radial distance from the origin, which refines the domain at radius $r=7M$ and once again at $r=3M$. As we adopt a three-partition [9], the volume size of each level are $[0.333M,0.111M,0.037M]$. Sommerfeld conditions (Section 3.5) suppress reflections along the domain boundary. No volume center coincides with the origin to avoid resolving numerical infinity at the puncture location.
For setup and running parameters, we adopt $\kappa_{1}=0.1,\kappa_{2}=0,\kappa_{3}=0.5,e=c=\tau=1.0$ and $\mu=0.2$ for a similar evolving system as the original CCZ4 system[11]. The fully functional gamma driver condition is employed for the evolution of the shift vector $\beta^{i}$ in the black hole system, and we set the parameters as $f=0.75$ and $\eta=1$ in Eqs. (23) and (29). The KO coefficient and the CFL ratio remain at $8.0$ and $0.1$, respectively, and the simulation runs until a code time of $T=240$. All figures below utilize $x:=r/M$ as the unit of distance for simplicity.
@@
Challenges arising around AMR boundaries
@@
Aggressive AMR introduces its specific challenges: with AMR, the evolution of a second-order formulation with KO damping remains stable until a distortion appears around the resolution boundary ($r=3M$), where the fine grid transitions into a coarser resolution. There, $K$ and $\alpha$ show a discontinuity of the gradient which gets amplified over time. $\tilde{\gamma}_{11}$ also exhibits fluctuations around the resolution boundary. The pattern is more clearly shown in Figure 12, where the profiles for violation of the Hamiltonian constraint $H$ are plotted along the positive x-axis, at six timestamps. Any violation around the black hole located at the center of the domain gets damped. However, the fluctuations along the resolution boundary ($r=3M$) continue to grow and start to propagate inwards. The simulation is likely to become unstable when the fluctuations reach the puncture.
@@
Any adaptive mesh resolution transition on non-conformal meshes is known to trigger reflections of outgoing waves: We have hanging vertices and hanging faces, i.e. fine grid vertices with less than @$2^{d}@f$ adjacent cells on the finest mesh level and faces where the left and the right adjacent cell are of different size. Let a wave travel from a fine mesh patch through a face into an adjacent coarser one. High-frequency modes stemming from the fine-resolution domain cannot be represented on an adjacent coarser grid and hence are reflected back into the the direction they are coming from. For dynamic simulations as encountered for moving black holes, these spurious reflections typically do not cause major stability issues. For the stationary black hole, however, we observe that the reflection waves are caught within the fine resolution domain and eventually amplify each other over time. They accumulate. A stationary black hole instead can create standing error waves that grow larger over time.
ExaGRyPE’s out-of-the-box linear interpolation and restriction schemes are not powerful enough to handle our highly non-linear PDE system over a long simulation span if we encounter a stationary black hole resolved with AMR. Without AMR, the simulation is computationally unfeasible. However, as long as black holes move, the KO dissipation is sufficient to compensate for any spurious reflections.
@@
${\delta t}\leq C\frac{h}{p^{2}},$
where we have a polynomial degree of $p=4$ for the FD4 solver and $p=1$ for FV. To match the time step sizes, the FV patch size is $4^{2}$ times larger than its FD4 cousin.
The profile of the $\phi$ field of the two solvers shows a stronger smoothness of the FV scheme around the black hole (Figure 13, lower panel). As the FV solution overwrites FD4 in the center, while the FD4 solution determines the halo data of the FV solution, thesolutions appear to overlap. However, only the FV scheme evolves logically.
@@
While the coupling is an appealing concept to create smooth, stable solutions without the need to use an artificial KO damping, i.e. solely due to first-order numerical principles, several challenges remain: tests with ExaGRyPE suggest that a direct overwrite of the FD4 solution can sometimes yield spurious waves similar to the AMR reflections. ExaGRyPE hence offers averaging and lets the “validity” of the FV scheme fade in: Directly around the black hole, the FV scheme dominates. Further away, FD4 yields a valid solution to and the FV solution is not even computed. In-between, we have a buffer zone. Both solvers run and the eventual result is a weighted average between the two solutions. The averaging shifts from “all FD4” over “predominantly FD4” over to “predominantly FV”. While this seems to yield higher-quality data, the precise effects have yet to be studied. It is also an open question of how big the FV region around the black hole has to be chosen overall. If it is too small, the damping effect of the FV region does not cover the unstable domain and we suffer from instabilities again. If it is too large, the numerical dissipation of FV pollutes the result in the region we are interested in. The biggest challenge arising from the coupled approach, however, is the likely difficult load balancing, as almost all computing effort is spent on a few FV patches of a very high resolution. We aim to investigate and explore more the potential of this coupling scheme in future work.

### v3

@@ 4.2.1 Solver Configuration
Once we have a reasonably fine initial mesh, ExaHyPE starts to query the solver objects if a certain area should be refined. In this stage, the solvers do not yet have a solution at hand. Either we hard-code the initial refinement pattern, or we make the algorithm run through a decision logic which analyses “if we had a solution here, would we refine”. This implements a feature-based initial refinement. The present case studies all use hard-coded patterns, adopting static grids where the finest resolution covering the whole interested region. The refinement criterion can be either written into the C++ template or injected the implementation of the C++ code directly from Python via my_solver.set_implementation(refinement_criterion
@@
. The dynamic grid refinement during evolution is also available in ExaGRyPE and can tag volumes for refinement either based on coordinates, similar to the “Tracker-based AMR” in [49], or based on variables, similar to the “tagging on gradient” in [50]. This dynamic feature is not used for the current benchmarks, and a full investigation is left for future work.
@@ 5.1 Gauge Wave
$\hat{t}$ $=$ $t-\frac{A}{2k\pi}\cos[k\pi(x-t)],$ (26)
$\hat{x}$ $=$ $x+\frac{A}{2k\pi}\cos[k\pi(x-t)],\qquad\text{and}$ (27)
$\hat{y}$ $=$ $y,\ \hat{z}=z,$ (28)
@@
$\alpha=\sqrt{H},~{}\beta^{i}=0,~{}\phi=H^{-1/6},\quad\text{and}$ (31)
$K_{xx}$ $=$ $-\frac{\partial_{t}H}{2\alpha}$ (32)
@@
$K_{ij,{\rm others}}$ $=$ $0,$ (33)
$K=\gamma^{ij}K_{ij}$ $=$ $\frac{K_{xx}}{H}$ (34)
@@
The first simulation test we report in this subsection utilizes a computational domain $\Omega=[-0.5,0.5]^{3}$ on a regular grid subject to periodic boundary conditions. Periodic boundary conditions are also a newly added feature released with ExaHyPE 2 as opposed to the first-generation code. We divide the domain into 162 volumes per dimension, thus leading to a volume size of $0.006173$. AMR is not enabled for this test.
@@
We illustrate the simulation result at code time $T=0.5$ in Figure 8. It shows the profiles of $\tilde{\gamma}_{11}$ along the $x$-axis for different kernels (FV, FD4-RK1, FD4-RK2 and FD4-RK4), employed with the first-order formulation. All the solvers yield the correct traveling speed and phase of the wave. However, the Finite Volume solver suffers from severe amplitude damping, and a similar dissipation is also observed in the black hole benchmarks. Consequently, we have chosen to use the FD4 scheme as our default solver.
@@
Figure 8: The profile of $\tilde{\gamma}_{11}$ along $x$-axis at code time $0.5$ for different solvers: FV-RK1 (red solid), FD4-RK1 (blue solid), FD4-RK2 (green dashed) and FD4-RK4 (orange dotted).
Figure 9: Convergence behavior of FD4 with three different time integrators for different CFL constants. The convergence of the final error depends on the relative magnitude of the contributing terms, i.e., spatial and temporal discretization errors. See the text for detailed explanations.
On the other hand, different orders of the RK schemes give rather consistent results against the analytical solution. To further investigate different effects from the temporal discretization scheme, we calculate the $L_{2}$ error of the Hamilton constraint as
$L_{2,H}:=\sqrt{\frac{\sum_{i}(H^{i}_{\rm simulated}-H_{\rm theoretical})^{2}}{n}},$ (35)
where $n$ is the total number of grid points, and we always have $H_{\rm theoretical}=0$. We conduct a convergence study for different orders of the RK scheme based on this accuracy metric at code time $t=0.4$ and report the result in Figure 9. A similar study can be conducted for any other reasonably small time stamp, and long-term effects will be the subject of study later. All physical and running parameters of tests here are identical to the first benchmark above except for the KO coefficient, which is set to 0.1.
In our explicit time-stepping scheme, three terms contribute to the $L_{2}$ error:
  1. 1.
    The spatial discretization error of the physical quantities, which is the same for all setups and scales with the mesh size in a convergence speed $\sim h^{4}$.
  2. 2.
    The time discretization error due to the Runge-Kutta scheme of choice. Through Eq. (20), the chosen time step size correlates linearly to the mesh size, thus following the convergence line $\sim h^{k}$ where $k$ is the order of the RK scheme.
  3. 3.
    The influence of the KO term, which is an artificial construct and hence constitutes an error in itself. It has a convergence speed $\sim h^{5}$ (see D).
How the $L_{2}$ error in our tests converges depends on which term takes the lead in magnitude. As shown in Figure 9, for the RK1 scheme, the temporal discretization error is quite large and results in an expected convergence speed $\sim h^{1}$ in the Hamiltonian constraint violation (blue lines with circle markers). RK2 and RK4 have relatively small temporal errors and if we have a too-large KO coefficient, then term 3 would overtake the leading term in error and yield an incorrect fifth-order convergence speed (not shown). When a proper (small) KO coefficient is set, the RK4 prediction also shows an expected fourth-order convergence (blue line with triangle markers). However, if the orders of spatial and temporal schemes do not match, the convergence of the final error will depend on the comparison of the magnitudes of each term. To demonstrate this, we plot two lines of the RK2 scheme with different ${\rm CFL}$ ratios, i.e., different sizes of the time steps. When the timestep is small (${\rm CFL}=0.1$, red dotted line with square markers), the RK2 scheme shows nearly identical error magnitude and convergence speed with the RK4 line, indicating we are in a regime dominated by spatial discretization errors. On the other hand, in the big timestep case (${\rm CFL}=1.5$, red solid line with square markers), the temporal discretization error is big enough and takes the lead, thus the line shows $\sim h^{2}$ convergence speed as we expected. We also identify more complex convergence behavior for the RK2 scheme when an intermediate time step size is used ($1.50.1<{\rm CFL}<$ 1.5). We hope to investigate the convergence properties of ExaGRyPE’s numerical schemes in more detail in future work.
Figure 10: Long-term evolution of the $L_{2}$ error with first-order (RK1) and second-order (RK2) time-integrator for the same gauge wave setup. The CFL constraint and KO term weight are chosen to dampen the constraint violation effectively. The low-order temporal scheme exhibits significant error accumulation, eventually leading to a crash, while the second-order scheme maintains a more stable error evolution throughout the simulation.
For a code with known convergence behavior and a reasonably chosen damping through the KO term and the CFL factor, we expect to be able to run long-term simulations. Simulations without a damping term are in turn expected to become unstable after a while [16]. This hypothesis is subject to our final gauge wave experiment. Unfortunately, our data suggest that a mere application of the KO correction plus a reasonably small CFL factor is not sufficient to construct a long-term stable solution (Figure 10): Even though the error seems to plateau after a while, it “suddenly” explodes within the plateauing region. Plots of the wave profile just before the error explosion reveal that the wave the pattern becomes skewed (not shown): The original sinusoidal pattern approaches a smoothed-out sawtooth profile, i.e. we observe some stiffening of the numerical solution. This is still a smooth effect, i.e. difficult for an artificial KO term to flag.
Convergence studies at a certain sample point consequently are insufficient to make statements on the long-term applicability of a code base. They might suggest that low-order time integrators are capable of long-term evolution of the system as long as we are willing to accept the resulting temporal discretization error. However, in practice, errors in the time evolution accumulate and eventually render the simulation unstable, i.e. make the scheme consistent but not converging. We need to employ a high-order temporal discretization scheme (RK2 or RK4) that matches our high spatial discretization order to construct stable schemes for the gauge wave. Our results demonstrate that a second-order temporal scheme has significantly reduced the $L_{2}$ error and stabilizes the simulation, as shown by the black lines in Figure 10.
The exact interplay of the KO term, CFL condition, Hamiltonian constraint violation, and long-term stability has to be the subject of future studies. We notably cannot make statements yet if this behavior results from an error accumulation that only arises for particular stationary or quasi-stationary solutions. Empirical data suggest that this is the case and that notably outflow conditions allow the error to escape from the domain (see the single black hole benchmark below). From hereon we pick reasonably small CFL factors and assume that the solution remains stable and consistent, while the arising error is dominated by the temporal discretization order.
@@ 5.2 Single Schwarzschild Black Hole
We next examine a single Schwarzschild black hole. This is yet another stationary test, where all numerical ingredients have to be carefully balanced against each other such that the initial setup effectively yields a long-term stable, time-invariant solution. The black hole is placed at the origin of the three-dimensional domain, with an ADM mass $M=1$ and zero spin $S=0$. All initial conditions for this test stem from the TwoPuncture module of the Einsteintoolkit library and the initial lapse is set to
$\alpha=\frac{1}{2}\left(\frac{1-\frac{1}{2}\left(M/r\right)}{1+\frac{1}{2}\left(M/r\right)}+1\right).$ (36)
Other initial profiles and lapses (e.g., a power of conformal factor) are also used in other numerical codes. However, convergence from different setups has been observed: different initial lapses evolve into nearly the same static solution. The initial shift vector is set to zero. Also, the extrinsic curvature $K_{ij}$ vanishes in the initial condition: its trace is zero as we assume the maximal slicing $K=0$, and its residual is zero as no linear or angular momentum is presented in this system. Finally, we also assume conformal flatness in the initial condition, i.e. $\tilde{\gamma}_{ij}=1$.
@@
$\phi\equiv\psi^{-2}=\left(1+\frac{M}{2r}\right)^{-2}.$ (37)
The domain in this test has a size of $[-9M,9M]^{3}$. We use three levels of SMR (Static Mesh Refinement) based on the radial distance from the origin, which refines the domain at radius $r=7M$ and once again at $r=3M$. As we adopt a three-partition [14], the volume size of each level are $[0.333M,0.111M,0.037M]$. Sommerfeld conditions (Section 3.5) suppress reflections along the domain boundary. No volume center coincides with the origin to avoid resolving numerical infinity at the puncture location.
For setup and running parameters, we adopt $\kappa_{1}=0.1,\kappa_{2}=0,\kappa_{3}=0.5,e=c=\tau=1.0$ and $\mu=0.2$ for a similar evolving system as the original CCZ4 system[16]. The fully functional gamma driver condition is employed for the evolution of the shift vector $\beta^{i}$ in the black hole system, and we set the parameters as $f=0.75$ and $\eta=1$ in Eqs. (53) and (59). The KO coefficient and the CFL ratio remain at $8.0$ and $0.1$, respectively, and the simulation runs until a code time of $T=240$ with the RK1 temporal scheme. All figures below utilize $x:=r/M$ as the unit of distance for simplicity.
@@
Challenges arising around refinement boundaries
@@
Aggressive SMR introduces its specific challenges: with SMR, the evolution of a second-order formulation with KO damping remains stable until a distortion appears around the resolution boundary ($r=3M$), where the fine grid transitions into a coarser resolution. There, $K$ and $\alpha$ show a discontinuity of the gradient which gets amplified over time. $\tilde{\gamma}_{11}$ also exhibits fluctuations around the resolution boundary. The pattern is more clearly shown in Figure 13, where the profiles for violation of the Hamiltonian constraint $H$ are plotted along the positive x-axis, at six timestamps. Any violation around the black hole located at the center of the domain gets damped. However, the fluctuations along the resolution boundary ($r=3M$) continue to grow and start to propagate inwards. The simulation is likely to become unstable when the fluctuations reach the puncture. This is confirmed by Figure 14, where we plot the evolution of the $L_{2}$ error of the Hamiltonian constraint up to code time $120$. The $L_{2}$ error in the first-order formulation increases quickly, departs from the stable second-order counterpart around 20 code time, and leads to a crash later. In contrast, the second-order formulation can provide a stable solution for the puncture all the time and shows continuous damping of the constraint violation until the errors at the refinement boundary start to take the lead and amplify the violations.
Figure 14: Evolution of the $L_{2}$ error of the Hamiltonian constraint $H$ and the magnitude of Momentum constraint $|M|$ for the single black hole scenario in two formulations. First-order and second-order formulations give nearly identical evolution initially, but eventually the former becomes unstable and crashes. The second-order formulation can run stably until the errors raised on the refinement boundary become non-negligible.
@@
Any adaptive mesh resolution transition on non-conformal meshes is known to trigger reflections of outgoing waves: We have hanging vertices and hanging faces, i.e. fine grid vertices with less than $2^{d}$ adjacent cells on the finest mesh level and faces where the left and the right adjacent cell are of different size. Let a wave travel from a fine mesh patch through a face into an adjacent coarser one. High-frequency modes stemming from the fine-resolution domain cannot be represented on an adjacent coarser grid and hence are reflected back into the the direction they are coming from. For dynamic simulations as encountered for moving black holes, these spurious reflections typically do not cause major stability issues. For the stationary black hole, however, we observe that the reflection waves are caught within the fine resolution domain and eventually amplify each other over time. They accumulate. A stationary black hole instead can create standing error waves that grow larger over time.
ExaGRyPE’s out-of-the-box linear interpolation and restriction schemes are not powerful enough to handle our highly non-linear PDE system over a long simulation span if we encounter a stationary black hole resolved with mesh refinement. However, as long as black holes move, the KO dissipation is sufficient to compensate for any spurious reflections. A more general solution is to implement higher-order interpolation and restriction schemes, which can provide higher accuracy around the refinement boundary. We have already observed a significant improvement in reflection suppression when we apply a second-order scheme in our recent tests of single black hole spacetimes (not shown). The refinement transition strategies play an important role in achieving high-quality numerical relativity simulations (e.g., see discussion in [49] and [50]), and we plan to report more comprehensively on how different strategies affect the simulation behavior of ExaGRyPE in future work.
@@
${\delta t}\leq C\frac{h}{p^{2}},$ (38)
where we have a polynomial degree of $p=4$ for the FD4 solver and $p=1$ for FV. To match the time step sizes, the FV patch size is $4^{2}$ times larger (per dimension) than its FD4 cousin, i.e., contains $16^{3}$ times more volumes per patch.
The profile of the $\phi$ field of the two solvers shows a stronger smoothness of the FV scheme around the black hole (Figure 15, lower panel). As the FV solution overwrites FD4 in the center, while the FD4 solution determines the halo data of the FV solution, the solutions appear to overlap. However, only the FV scheme evolves without interference.
@@
While the coupling is an appealing concept to create smooth, stable solutions without the need to use an artificial KO damping, i.e. solely due to first-order numerical principles, several challenges remain: tests with ExaGRyPE suggest that a direct overwrite of the FD4 solution can sometimes yield spurious waves similar to the reflections from the refinement boundary. ExaGRyPE hence offers averaging and lets the “validity” of the FV scheme fade in: Directly around the black hole, the FV scheme dominates. Further away, FD4 yields a valid solution to and the FV solution is not even computed. In-between, we have a buffer zone. Both solvers run and the eventual result is a weighted average between the two solutions. The averaging shifts from “all FD4” over “predominantly FD4” over to “predominantly FV”. While this seems to yield higher-quality data, the precise effects have yet to be studied. It is also an open question of how big the FV region around the black hole has to be chosen overall. If it is too small, the damping effect of the FV region does not cover the unstable domain and we suffer from instabilities again. If it is too large, the numerical dissipation of FV pollutes the result in the region we are interested in. The biggest challenge arising from the coupled approach, however, is the likely difficult load balancing, as almost all computing effort is spent on a few FV patches of a very high resolution. We aim to investigate and explore more the potential of this coupling scheme in future work.